Solve equations by factoring
Solve the equation.
x^3 + 13x^2 + 40x = 0

Answer by ewatrrr(24785)


Hi
Solving equation by factoring
x^3 + 13x^2 + 40x = 0 Note ALL terms have x in common
x(x^2 + 13x + 40) = 0
x(x+8)(x+5) = 0 Note:SUM of the inner product(8x) and the outer product(5x) = 13x
x = 0
(x+8)= 0
x = -8
(x+5) = 0
x = -5